# Curvature-Constrained Vector Field for Motion Planning of Nonholonomic Robots

### v1

@@ Appendix E Proof of Theorem 3
$\mathcal{S}$ =$=\left\{\bm{g}\mid r_{\delta}<\rho\right\}=\left\{\bm{g}\mid k(r_{\delta})<A(r_{\delta})\right\}$ (46)
$\supseteq\left\{\bm{g}\mid\bar{\omega}_{0}>\bar{\omega}\right\}\supseteq\left\{\bm{g}\mid\left|\omega_{0}\right|>\bar{\omega}\right\}$
@@ Appendix F Proof of Theorem 4
which means the feedback term $-k_{\omega}\theta_{e}$ and the feedforward term $\omega_{r}$ in (20) have different signs. In addition, the dynamic gain (23) ensures that $\lvert k_{\omega}\theta_{e}\rvert\leq\bar{\omega}$. Therefore, we can conclude that (22) saturates, i.e. $\lvert\omega_{0}\rvert=\lvert-k_{\omega}\theta_{e}+\omega_{r}\rvert>\bar{\omega}$, if and only if $\lvert\omega_{r}\rvert>\bar{\omega}+\lvert k_{\omega}\theta_{e}\rvert$. When $\theta_{e}>0$ and the angular velocity saturates, there holds that
@@ Appendix G Proof of Theorem 5
$\dot{V}_{\delta}<0,\ \forall\lvert r_{\delta}-r_{2}\rvert\geq\gamma(\lvert\theta_{e}\rvert),$
where $\gamma(\lvert\theta_{e}\rvert)=\max{\left\{\gamma_{1},\gamma_{2}\right\}}$, $\gamma_{1}(\lvert\theta_{e}\rvert)=r_{2}-h^{-1}(\pi/2-\lvert\theta_{e}\rvert))\leq r_{2}-r_{1}$ and $\gamma_{2}(\lvert\theta_{e}\rvert)=h^{-1}(\pi/2+\lvert\theta_{e}\rvert))-r_{2}\leq r_{3}-r_{2}$ are class $\mathcal{K}$ functions as illustrated in Figure 14. Referring to [44, Theorem 4.19], the above Lyapunov analysis suggests that the subsystem (48a) with bounded input $\lvert\theta_{e}\rvert\leq\pi/2$ is input-to-state stable (ISS). Furthermore, since Theorem 4 indicates that the origin of (48b) is asymptotically stable, we conclude the cascade system (48) converges to the equilibrium where $r_{\delta}=r_{2}$, $\theta_{e}=0$ [44, Lemma 4.7].
@@ Acknowledgment
The authors would like to thank Xiwang Dong, Haoze Dong and Tianqing Liu for the implementation of the experiments with the Ackermann UGV and the HIL experiments with fixed-wing UAV.

### v2

@@ Appendix E Proof of Theorem 3
$\mathcal{S}$ $=\left\{\bm{g}\mid r_{\delta}<\rho\right\}=\left\{\bm{g}\mid k(r_{\delta})<A(r_{\delta})\right\}$ (48)
$\supseteq\left\{\bm{g}\mid\bar{\omega}_{0}>\bar{\omega}\right\}\supseteq\left\{\bm{g}\mid\left|\omega_{0}\right|>\bar{\omega}\right\},$
@@ Appendix F Proof of Theorem 4
which means the feedback term $k_{\omega}\theta_{e}$ and the feedforward term $\omega_{r}$ in (20) have different signs. In addition, the dynamic gain (23) ensures that $\lvert k_{\omega}\theta_{e}\rvert\leq\bar{\omega}$. Therefore, we can conclude that (22) saturates, i.e. $\lvert\omega_{0}\rvert=\lvert-k_{\omega}\theta_{e}+\omega_{r}\rvert>\bar{\omega}$, if and only if $\lvert\omega_{r}\rvert>\bar{\omega}+\lvert k_{\omega}\theta_{e}\rvert$. When $\theta_{e}>0$ and the angular velocity saturates, there holds that
@@ Appendix G Proof of Theorem 5
[…]
where $\gamma(\lvert\theta_{e}\rvert)=\max{\left\{\gamma_{1},\gamma_{2}\right\}}$, $\gamma_{1}(\lvert\theta_{e}\rvert)=r_{2}-h^{-1}(\pi/2-\lvert\theta_{e}\rvert))\leq r_{2}-r_{1}$ and $\gamma_{2}(\lvert\theta_{e}\rvert)=h^{-1}(\pi/2+\lvert\theta_{e}\rvert))-r_{2}\leq r_{3}-r_{2}$ are class $\mathcal{K}$ functions as illustrated in Figure 16. Referring to [khalil2002nonlinear, Theorem 4.19], the above Lyapunov analysis suggests that the subsystem (50a) with bounded input $\lvert\theta_{e}\rvert\leq\pi/2$ is input-to-state stable (ISS). Furthermore, since Theorem 4 indicates that the origin of (50b) is asymptotically stable, we conclude the cascade system (50) converges to the equilibrium where $r_{\delta}=r_{2}$, $\theta_{e}=0$ [khalil2002nonlinear, Lemma 4.7].
@@ Acknowledgment
The authors would like to thank Xiwang Dong, Haoze Dong, and Tianqing Liu for their assistance with the implementation of the experiments on the Ackermann UGV and the fixed-wing UAV.
[Uncaptioned image] Yike Qiao (Student Member, IEEE) is working toward the Ph.D. degree in dynamical systems and control with the School of Advanced Manufacturing and Robotics, Peking University, Beijing, China. His research interests include motion planning and control of mobile robots under kinematic constraints.
[Uncaptioned image] Xiaodong He (Member, IEEE) received the B.Eng. degree in mechanical engineering from China University of Petroleum, Qingdao, China, in 2016, and the Ph.D. degree in dynamics and control from the Peking University, Beijing, China, in 2021. From 2021 to 2024, he was a Postdoctoral Researcher and an Associate Researcher with the College of Engineering, Peking University. Since 2024, he has been an Associate Professor with the School of Automation and Electrical Engineering, University of Science and Technology Beijing, Beijing, China. His research interests include motion planning and control of nonholonomic systems. Dr. He was a finalist for the Zhang Si-Ying Outstanding Youth Paper Award of Chinese Control and Decision Conference (CCDC) in 2023.
[Uncaptioned image] An Zhuo is a Ph.D. candidate at the School of Advanced Manufacturing and Robotics, Peking University. He received his bachelor’s degree from the National University of Defense Technology in 2024. His research interests include unmanned swarm systems, cooperative control, and decision making.
[Uncaptioned image] Zhiyong Sun (Member, IEEE) received the Ph.D. degree from The Australian National University (ANU), Canberra ACT, Australia, in February 2017. He was a Research Fellow/Lecturer with the Research School of Engineering, ANU, from 2017 to 2018. From June 2018 to January 2020, he worked as a postdoctoral researcher at the Department of Automatic Control, Lund University of Sweden. From January 2020 to July 2024, he worked with Eindhoven University of Technology (TU/e) as an assistant professor. He joined Peking University of China as a faculty member in the summer of 2024. His research interests include multi-agent systems, control of autonomous formations, distributed control and optimization. He has authored the book titled Cooperative Coordination and Formation Control for Multiagent Systems (Springer, 2018). He has won the Springer Best Ph.D. Thesis Award, and several best paper and student paper awards from IEEE CDC, AuCC, ICRA, CCTA and ICCA.
[Uncaptioned image] Weimin Bao received the B.E. degree in electronic engineering from Xidian University, Xi’an, China, in 1982. He is currently a CAS Member of China and a Professor with Xidian University. His research interests include space launch vehicle and control engineering.
[Uncaptioned image] Zhongkui Li (Senior Member, IEEE) received the B.S. degree in space engineering from the National University of Defense Technology, China, in 2005, and his Ph.D. degree in dynamical systems and control from Peking University, China, in 2010. Since 2013, Dr. Li has been with the Peking University, China, where he is currently a Full Professor with the School of Advanced Manufacturing and Robotics. His current research interests include cooperative control and planning of multi-agent systems. Dr. Li was the recipient of the China National Science Funds for Distinguished Young Scientists in 2024, the State Natural Science Award of China in 2015, the Natural Science Award of the Ministry of Education of China in 2022 and 2011, and the National Excellent Doctoral Thesis Award of China in 2012. His coauthored papers received the IET Control Theory & Applications Premium Award in 2013 and the Best Paper Award of Journal of Systems Science & Complexity in 2012. He serves/has served as an Associate Editor of IEEE Transactions on Automatic Control, and several other journals.